OpenStudy (gabylovesyou):

use the table to answer this question: x y 1 10 3 4 5 -2 Find the average rate of change for the given function from x = 3 to x = 5. -6 -3 3 6

OpenStudy (gabylovesyou):

@ganeshie8

ganeshie8 (ganeshie8):

simply find the slope

ganeshie8 (ganeshie8):

x y 3 4 5 -2

OpenStudy (gabylovesyou):

slope = y2 - y1 / x2 - x1 right ?

ganeshie8 (ganeshie8):

yes

OpenStudy (gabylovesyou):

-2 - 4 = -6 5 - 3 = 2 -3 is the slope

ganeshie8 (ganeshie8):

\[\huge \color{red}{\checkmark}\]

OpenStudy (gabylovesyou):

ty